OpenStudy (anonymous):

what would be the area of a twelve inch pizza? im trying to find it but cant get the right answer

OpenStudy (anonymous):

12 inch = diameter ?

OpenStudy (anonymous):

thats the problem

OpenStudy (anonymous):

i think so but i dont know. the question doesnt say

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

if that's the case then: \[A = \pi \times (\frac{ diameter }{2 })^2\]where pi=3.14 diameter/2 = 6 inch

OpenStudy (anonymous):

i know

OpenStudy (anonymous):

oh wait i found it

OpenStudy (anonymous):

its 110

OpenStudy (anonymous):

yeah is that a possible answer ?

OpenStudy (anonymous):

YES